OpenStudy (anonymous):

Solve the equation. 3x^2+96=0

OpenStudy (noseboy908):

first do what you can do first to isolate the variable

OpenStudy (noseboy908):

get 96 on it's own side of the equals

OpenStudy (anonymous):

96 +-96 +3x^2=0 +-96

OpenStudy (noseboy908):

pretty much, so what do you have now?

OpenStudy (anonymous):

okay so this is what I have so far...\[3(x ^{2}+32)=0\] =\[3\sqrt{x ^{2}} = \sqrt{32}\] =\[x=\sqrt{8}\sqrt{4}\] and I kept going and got x=4, but my teacher said it's wrong so if you could explain to me what I did wron that would be great.

OpenStudy (noseboy908):

you are REALLY over complicating this, just tell me what you have after the first step

OpenStudy (anonymous):

3x^2=-96

OpenStudy (noseboy908):

ok, so what would you do next to isolate x?

OpenStudy (anonymous):

dive 3 to both sides?

OpenStudy (noseboy908):

yep

OpenStudy (noseboy908):

to get what?

OpenStudy (anonymous):

x^2=-32 & now I would Square root both sides correct?

OpenStudy (noseboy908):

yes, to get?

OpenStudy (anonymous):

\[x=2i\sqrt{2}\]

OpenStudy (noseboy908):

that would appear correct if you simplified correctly. IF so, you're done.

OpenStudy (anonymous):

wait I think I simplified it wrong...

OpenStudy (anonymous):

\[x=2i\sqrt{2}\]